### v3

# Negative and positive anisotropic thermal expansion in 2D fullerene networks

Abstract
We find a design principle for tailoring thermal expansion properties in nearly-spherical molecular networks. Using 2D fullerene networks as a representative system, we realize positive thermal expansion along intermolecular [2 + 2] cycloaddition bonds and negative thermal expansion along intermolecular CC single bonds by varying the structural frameworks of molecules. The microscopic mechanism originates from a combination of the framework’s geometric flexibility and its transverse vibrational characteristics. Based on this insight, we find molecular networks beyond C60 with tunable thermal expansion. These findings shed light on the fundamental mechanisms governing thermal expansion in molecular networks towards rational materials design.
Thermal expansion is a fundamental property of materials that indicates increases in length, area, or volume upon heating, which is important in applications such as construction Johnson and Parsons (1944), seismographs Mohn (1999), and aerospace design Strock (1992); Toropova and Steeves (2015). Positive thermal expansion occurs as a result of an anharmonic potential energy surface, where interatomic distance increases with increasing temperature Barrera et al. (2005); Liu et al. (2017). Counterintuitively, some materials exhibit negative thermal expansion, where increasing temperature leads to a contraction along certain crystallographic directions Baughman and Galvão (1993); Mary et al. (1996); Das et al. (2010); Fortes et al. (2011); Takenaka (2012); Liu et al. (2018); Ritz and Benedek (2018); Shi et al. (2021); Li et al. (2022). Such behaviors are attributed to flexible crystalline networks Goodwin et al. (2008); Dove and Fang (2016), rigid unit modes Pryde et al. (1996); Sleight (1998); Tucker et al. (2005); Tan et al. (2024), and transverse displacements of bridging atoms Hancock et al. (2004); Goodwin and Kepert (2005) or membranes Huang et al. (2016a); Koocher et al. (2021). However, a general design principle for developing materials with negative thermal expansion is still lacking.
Recent synthesis of monolayer C60 networks Hou et al. (2022) provides new avenues for designing materials with tunable thermal expansion. These networks exhibit diverse crystalline frameworks Hou et al. (2022); Peng (2022, 2023); Jones and Peng (2023); Shearsby et al. (2025); Kayley and Peng (2025); Peng and Pizzochero (2025a, b); Meirzadeh et al. (2023) with nearly-spherical, stable units Tromer et al. (2022); Ribeiro et al. (2022) beyond rigid unit modes, as well as various intermolecular bridge bonds Zhao et al. (2024) with tunable transverse displacements. Compared to thermal expansion in C60 molecules and solids Nagel et al. (1999); Arvanitidis et al. (2003); Kwon et al. (2004); Brown et al. (2006), the thermal behavior of C60 monolayers has yet to be understood. In 2D form, the rotational degree of freedom of C60 leads to different types of intermolecular bonds in varied crystalline networks [Fig. 1(a)]. In this context, it would be insightful to study whether thermal expansion in molecular networks can be controlled by intermolecular bridge bonds.
Here we show that intermolecular bonds govern thermal expansion behaviors in fullerene networks. Interfullerene [2 + 2] cycloaddition bonds yield positive thermal expansion, whereas thermal contraction is found along single bonds. We identify the microscopic mechanism by analyzing the geometric flexibility of these bridge bonds. Remarkably, we find that the low-frequency transverse vibrations along the single bonds favor thermal contraction, in contrast to the transverse vibrations associated with the [2 + 2] cycloaddition bonds. Based on this understanding, we rationally design molecular networks beyond C60 with tailored thermal expansion.
Thermal expansion is simulated under the quasi-harmonic approximation Huang et al. (2016b); Pallikara and Skelton (2021) with volume-dependent phonons computed from density functional perturbation theory Gonze (1995); Baroni et al. (2001) using VASP Kresse and Furthmüller (1996a, b). The Gibbs free energy is obtained by finding the unique minimum value of the Helmholtz free energy Dove (1993); Togo et al. (2008); Togo and Tanaka (2015) with varied lattice constants and at a strain step of 0.2% Peng et al. (2019); Peng (2023).
Figure 1(a) shows the crystal structures of two distinct networks of C60 monolayers, namely, the quasi-tetragonal phase (qTP) and quasi-hexagonal phase (qHP). For qTP, each carbon cage is connected by vertical and horizontal [2+2] cycloaddition bonds along the and directions, respectively. Such bonds are expected to expand rigidly along their axis upon heating. In qHP, only the buckyballs along are linked by the [2+2] cycloaddition bonds, while CC single bonds link the neighboring cages along . Figure 1(b) shows the thermal expansion of the two phases. While qTP along both directions has positive thermal expansion, negative thermal expansion along is found for qHP up to 500 K, in contrast to the positive thermal expansion along (here we neglect the tiny negative thermal expansion along which might come from the soft vibrational modes of the isolated molecule with ellipsoidal deformation that keeps the surface area constant Kwon et al. (2004); Brown et al. (2006)).
To understand the thermal expansion behaviors, we study the geometric flexibility of the two phases. Figure 2 shows the variations of structural parameters for qTP and qHP monolayer networks at varied strains along and . These structural parameters measure the geometric flexibility of both intermolecular bonds () and individual molecules () along and , as shown in Fig. 2(a) and (c) for qTP and qHP, respectively.
For qTP, both the intermolecular bond and the molecular width expand rigidly upon uniaxial strains along . However, and remain nearly unchanged for strains along the other direction , as shown by the color map in Fig. 2(b). The same conclusion also holds for and .
For qHP, the intermolecular [2 + 2] cycloaddition bonds and molecular width also expand rigidly for parallel strains along , while resisting deformations for perpendicular strains along , as shown in Fig. 2(d), exhibiting behaviors similar to the [2 + 2] cycloaddition bonds in qTP. However, the intermolecular single bonds and molecular width in qHP respond differently to the strains. As shown by the color map in Fig. 2(d), the in qHP C60 becomes shorter with increased strain along . This notable contraction along with increasing indicates that the single bonds deform more readily, with compression specifically along being favorable for positive strains along . This hinge-like motion is expected in the single bonds as they are less resistant to perpendicular strains. On the other hand, the in qHP C60 expands upon strains along . Therefore, we can attribute negative thermal expansion in qHP to the geometric flexibility of the intermolecular single bonds instead of the molecules themselves.
The overall behavior demonstrates general features for different types of intermolecular bonds. The [2 + 2] cycloaddition bonds expand rigidly when strain is applied parallel to their direction while resisting deformations against perpendicular strains. Therefore, qTP C60 networks exhibit nearly-isotropic positive thermal expansion. On the other hand, the single bonds allow for hinge-like compression when applying perpendicular strains. Unlike qTP, the flexible single bonds in qHP contract like hinges along when the rigid [2 + 2] cycloaddition bonds along expand upon heating. Our findings demonstrate a distinctive interplay between flexibility-driven lattice contraction via CC single bonds and rigidity-induced structural expansion through [2 + 2] cycloaddition bonds. This drives a strong anisotropic thermal response in qHP C60, as summarized by the schematics in Fig. 1(a).
To further explore the microscopic mechanism of negative and positive anisotropic thermal expansion in qHP fullerene networks, we examine their vibrational modes. At 300 K, only phonons below 0.6 (4.3) THz have an occupation number above 10 (1) according to the Bose-Einstein distribution. Thus, we focus on low-frequency phonons hereafter. Figure 3 shows the phonon dispersion curves and the corresponding mode Grüneisen parameters . The phonon dispersion along Y– corresponds to the vibrations associated with the intermolecular [2 + 2] cycloaddition bonds along . The mode Grüneisen parameters of these vibrations are either positive or near zero, leading to thermal expansion. In contrast, the transverse displacements associated with the intermolecular single bonds have large negative along –X. There are two transverse phonon branches with the largest negative : an out-of-plane acoustic mode (ZA) with the coherent movement of all molecules along , and an out-of-plane optical mode (ZO) with alternating displacements between neighboring molecules, as illustrated by the schematics in Fig. 3. Both modes with large negative favor lattice contraction. In comparison, the longitudinal acoustic mode (LA) has positive , which contributes to lattice expansion instead but less strongly () than the ZA and ZO modes (). The overall picture confirms that transverse displacements of the single bonds yield thermal contraction, while the [2 + 2] cycloaddition bond oscillations contribute to expansion behaviors.
The rigidity of the [2 + 2] cycloaddition bonds, as well as the flexibility of the CC single bonds, provide a universal design principle to tailor thermal expansion behaviors. The [2 + 2] cycloaddition bonds between molecular cages impose structural, elastic, and vibrational constraints that only allow positive thermal expansion. Contrastingly, the flexible single bonds bridging the molecules allow for lattice contraction when perpendicular strains are applied, and the vibrational modes associated with these single bonds yield strong transverse displacements that favor thermal contraction. The flexibility of the single bonds is therefore the main driving factor in negative thermal expansion. Similar thermal behaviors in fullerene-based networks have also been found in previous molecular dynamics simulations Mortazavi (2023). Intuitively, we can either realize positive thermal expansion through intermolecular [2 + 2] cycloaddition bonds to resist lattice contraction, or utilize less rigid intermolecular single bonds to yield negative thermal expansion.
The discovery that CC single and [2 + 2] cycloaddition bonds contribute distinctively to thermal behavior provides a predictive tool for the rational design of thermally responsive materials with nearly-spherical building blocks such as icosahedral B12 units Kah et al. (2015); Sun et al. (2017); Zhang et al. (2017), fullerene cages Kroto (1987); Hallett et al. (1995), and all-metal clusters Xu et al. (2023) (for thermal expansion of 2D icosahedral B12 networks, see Fig. 5 in the End Matter). As a proof-of-principle study, we extend this design principle to monolayer qHP C24 networks, where the C24 molecules are linked through similar single bonds along but three intermolecular bonds between six carbon atoms along Wu and Peng (2025). The crystal structure of qHP C24 is shown in Fig. 4(a). Unlike the nearly-planar single bonds in qHP C60, the single bonds in C24 monolayers have a buckled structure owing to a larger molecular curvature and asymmetric intermolecular bonding positions. This, in combination with the smaller molecular size, leads to a higher density of interfullerene bonds and larger elastic constants than qHP C60 Wu and Peng (2025). It is therefore expected that the single bonds in C24 are more rigid than those in qHP C60, leading to stronger resistance to perpendicular deformations. Figure 4(b) shows smaller negative thermal expansion in C24 compared to C60 along as expected. Additionally, the positive thermal expansion along in C24 is also larger than that in C60. This is unsurprising since the three intermolecular bonds along in C24 are much more rigid than the [2 + 2] cycloaddition bonds in monolayer qHP C60 networks.
The recent experimental realization of monolayer fullerene networks Hou et al. (2022) offers a timely opportunity to directly test our predictions, which influence numerous fullerene-based applications such as device integration in photodetectors Zhang et al. (2025) and catalytic activity in photocatalysis Peng (2022); Jones and Peng (2023); Shearsby et al. (2025); Wang et al. (2023). The temperature-dependent structural properties can be measured using high-resolution capacitance dilatometers Nagel et al. (1999) and synchrotron X-ray diffraction Arvanitidis et al. (2003). The low-frequency phonon modes can be detected by Raman and infrared spectroscopy Brown et al. (2006), enabling verification of the predicted transverse vibrational mechanisms. From an application perspective, molecular networks with tunable thermal expansion present compelling possibilities. Materials with designed negative or near-zero thermal expansion are in high demand in flexible electronics Mohn (1999), aerospace composites Strock (1992), and precision instrumentation Wang et al. (2024), where thermal mismatch must be minimized. Our results indicate that such properties can be achieved not through complex multi-phase composites, but through intrinsic molecular architecture. Previous experimental data on thermal expansion in molecular networks Baughman and Galvão (1993); Goodwin et al. (2008); Fortes et al. (2011); Dove and Fang (2016); Liu et al. (2018); Yao et al. (2019); Bond (2021); van der Lee and Dumitrescu (2021); Saha and Veluthaparambath (2024) including polymeric fullerene Blank et al. (1998); Nagel et al. (1999) confirm the fact that thermal expansion in molecular crystals can be controlled by bonding motifs and crystalline geometries, and similar mechanisms have been observed experimentally such as flexibility-driven “hinge-like” motion in crystalline networks Baughman and Galvão (1993); Goodwin et al. (2008); Fortes et al. (2011); Dove and Fang (2016) and transverse displacements of bridging atoms Hancock et al. (2004); Goodwin and Kepert (2005) or membranes Huang et al. (2016a); Koocher et al. (2021). This, again, supports the broader generalisability of our proposed mechanism.
In summary, we establish a general method for engineering thermal expansion in cage-like molecular networks through different intermolecular bonds. Using fullerene as a representative system, we show that intermolecular [2 + 2] cycloaddition bonds result in positive thermal expansion, whereas CC single bonds lead to negative thermal expansion. By varying the structural frameworks of fullerene molecules, we can design a nearly-square lattice with positive thermal expansion through [2 + 2] cycloaddition bonds along both in-plane directions. Similarly, we can also realize negative and positive anisotropic thermal expansion in a nearly-triangular lattice through single bonds along one direction and [2 + 2] cycloaddition bonds along the other direction. The origin is related to both the geometric flexibility of different intermolecular bonds and their corresponding transverse vibrations. By identifying and uncovering this microscopic mechanism, we can further design molecular networks with tailored thermal expansion.
Acknowledgements.
A.S. acknowledges support from Homerton College Cambridge for a Homerton-Victoria Brahm Schild Grant. J.W. acknowledges support from the Cambridge Undergraduate Research Opportunities Programme and from Peterhouse for the James Porter Scholarship. B.P. acknowledges support from Magdalene College Cambridge for a Nevile Research Fellowship. The calculations were performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1), and DiRAC funding from the Science and Technology Facilities Council (http://www.dirac.ac.uk ), as well as with computational support from the UK Materials and Molecular Modelling Hub, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1), for which access was obtained via the UKCP consortium and funded by EPSRC grant ref EP/P022561/1.References
- Johnson and Parsons (1944) Walter H Johnson and Willard H Parsons, Thermal expansion of concrete aggregate materials (US Government Printing Office Washington, DC, USA, 1944).
- Mohn (1999) Peter Mohn, “A century of zero expansion,” Nature 400, 18–19 (1999).
- Strock (1992) John D. Strock, “Development of zero coefficient of thermal expansion composite tubes for stable space structures,” Proc. SPIE 1690, 223–230 (1992).
- Toropova and Steeves (2015) Marina M. Toropova and Craig A. Steeves, “Adaptive bimaterial lattices to mitigate thermal expansion mismatch stresses in satellite structures,” Acta Astronautica 113, 132–141 (2015).
- Barrera et al. (2005) G D Barrera, J A O Bruno, T H K Barron, and N L Allan, “Negative thermal expansion,” Journal of Physics: Condensed Matter 17, R217 (2005).
- Liu et al. (2017) Zi-Kui Liu, Shun-Li Shang, and Yi Wang, “Fundamentals of thermal expansion and thermal contraction,” Materials 10, 410 (2017).
- Baughman and Galvão (1993) Ray H. Baughman and Douglas S. Galvão, “Crystalline networks with unusual predicted mechanical and thermal properties,” Nature 365, 735–737 (1993).
- Mary et al. (1996) T. A. Mary, J. S. O. Evans, T. Vogt, and A. W. Sleight, “Negative thermal expansion from 0.3 to 1050 kelvin in zrw2o8,” Science 272, 90–92 (1996).
- Das et al. (2010) Dinabandhu Das, Tia Jacobs, and Leonard J. Barbour, “Exceptionally large positive and negative anisotropic thermal expansion of an organic crystalline material,” Nature Materials 9, 36–39 (2010).
- Fortes et al. (2011) A. Dominic Fortes, Emmanuelle Suard, and Kevin S. Knight, “Negative linear compressibility and massive anisotropic thermal expansion in methanol monohydrate,” Science 331, 742–746 (2011).
- Takenaka (2012) Koshi Takenaka, “Negative thermal expansion materials: technological key for control of thermal expansion,” Science and Technology of Advanced Materials 13, 013001 (2012).
- Liu et al. (2018) Zhanning Liu, Qilong Gao, Jun Chen, Jinxia Deng, Kun Lin, and Xianran Xing, “Negative thermal expansion in molecular materials,” Chem. Commun. 54, 5164–5176 (2018).
- Ritz and Benedek (2018) Ethan T. Ritz and Nicole A. Benedek, “Interplay between phonons and anisotropic elasticity drives negative thermal expansion in pbtio3,” Phys. Rev. Lett. 121, 255901 (2018).
- Shi et al. (2021) Naike Shi, Yuzhu Song, Xianran Xing, and Jun Chen, “Negative thermal expansion in framework structure materials,” Coordination Chemistry Reviews 449, 214204– (2021).
- Li et al. (2022) Qiang Li, Kun Lin, Zhanning Liu, Lei Hu, Yili Cao, Jun Chen, and Xianran Xing, “Chemical diversity for tailoring negative thermal expansion,” Chem. Rev. 122, 8438–8486 (2022).
- Goodwin et al. (2008) Andrew L. Goodwin, Mark Calleja, Michael J. Conterio, Martin T. Dove, John S. O. Evans, David A. Keen, Lars Peters, and Matthew G. Tucker, “Colossal positive and negative thermal expansion in the framework material ag3[co(cn)6],” Science 319, 794–797 (2008).
- Dove and Fang (2016) Martin T Dove and Hong Fang, “Negative thermal expansion and associated anomalous physical properties: review of the lattice dynamics theoretical foundation,” Reports on Progress in Physics 79, 066503 (2016).
- Pryde et al. (1996) Alexandra K A Pryde, Kenton D Hammonds, Martin T Dove, Volker Heine, Julian D Gale, and Michele C Warren, “Origin of the negative thermal expansion in ZrW2O8 and ZrV2O7,” Journal of Physics: Condensed Matter 8, 10973 (1996).
- Sleight (1998) A. W. Sleight, “Compounds that contract on heating,” Inorg. Chem. 37, 2854–2860 (1998).
- Tucker et al. (2005) Matthew G. Tucker, Andrew L. Goodwin, Martin T. Dove, David A. Keen, Stephen A. Wells, and John S. O. Evans, “Negative thermal expansion in zrw2o8: Mechanisms, rigid unit modes, and neutron total scattering,” Phys. Rev. Lett. 95, 255501 (2005).
- Tan et al. (2024) Lei Tan, Volker Heine, Gong Li, and Martin T Dove, “The rigid unit mode model: review of ideas and applications,” Reports on Progress in Physics 87, 126501 (2024).
- Hancock et al. (2004) Jason N. Hancock, Chandra Turpen, Zack Schlesinger, Glen R. Kowach, and Arthur P. Ramirez, “Unusual low-energy phonon dynamics in the negative thermal expansion compound zrw2o8,” Phys. Rev. Lett. 93, 225501 (2004).
- Goodwin and Kepert (2005) Andrew L. Goodwin and Cameron J. Kepert, “Negative thermal expansion and low-frequency modes in cyanide-bridged framework materials,” Phys. Rev. B 71, 140301 (2005).
- Huang et al. (2016a) Liang-Feng Huang, Xue-Zeng Lu, and James M. Rondinelli, “Tunable negative thermal expansion in layered perovskites from quasi-two-dimensional vibrations,” Phys. Rev. Lett. 117, 115901 (2016a).
- Koocher et al. (2021) Nathan Z. Koocher, Liang-Feng Huang, and James M. Rondinelli, “Negative thermal expansion in the ruddlesden-popper calcium titanates,” Phys. Rev. Mater. 5, 053601 (2021).
- Hou et al. (2022) Lingxiang Hou, Xueping Cui, Bo Guan, Shaozhi Wang, Ruian Li, Yunqi Liu, Daoben Zhu, and Jian Zheng, “Synthesis of a monolayer fullerene network,” Nature 606, 507–510 (2022).
- Peng (2022) Bo Peng, “Monolayer fullerene networks as photocatalysts for overall water splitting,” J. Am. Chem. Soc. 144, 19921–19931 (2022).
- Peng (2023) Bo Peng, “Stability and strength of monolayer polymeric c60,” Nano Lett. 23, 652–658 (2023).
- Jones and Peng (2023) Cory Jones and Bo Peng, “Boosting photocatalytic water splitting of polymeric c60 by reduced dimensionality from two-dimensional monolayer to one-dimensional chain,” J. Phys. Chem. Lett. 14, 11768–11773 (2023).
- Shearsby et al. (2025) Dylan Shearsby, Jiaqi Wu, Dekun Yang, and Bo Peng, “Tuning electronic and optical properties of 2d polymeric c60 by stacking two layers,” Nanoscale 17, 2616–2620 (2025).
- Kayley and Peng (2025) Darius Kayley and Bo Peng, “C60 building blocks with tuneable structures for tailored functionalities,” Computational Materials Today 6, 100030 (2025).
- Peng and Pizzochero (2025a) Bo Peng and Michele Pizzochero, “Monolayer C60 networks: a first-principles perspective,” Chem. Commun. 61, 10287–10302 (2025a).
- Peng and Pizzochero (2025b) Bo Peng and Michele Pizzochero, “Electronic structure of fullerene nanoribbons,” ACS Nano (2025b), 10.1021/acsnano.5c08991.
- Meirzadeh et al. (2023) Elena Meirzadeh, Austin M. Evans, Mehdi Rezaee, Milena Milich, Connor J. Dionne, Thomas P. Darlington, Si Tong Bao, Amymarie K. Bartholomew, Taketo Handa, Daniel J. Rizzo, Ren A. Wiscons, Mahniz Reza, Amirali Zangiabadi, Natalie Fardian-Melamed, Andrew C. Crowther, P. James Schuck, D. N. Basov, Xiaoyang Zhu, Ashutosh Giri, Patrick E. Hopkins, Philip Kim, Michael L. Steigerwald, Jingjing Yang, Colin Nuckolls, and Xavier Roy, “A few-layer covalent network of fullerenes,” Nature 613, 71–76 (2023).
- Tromer et al. (2022) Raphael M. Tromer, Luiz A. Ribeiro, and Douglas S. Galvão, “A DFT study of the electronic, optical, and mechanical properties of a recently synthesized monolayer fullerene network,” Chemical Physics Letters 804, 139925 (2022).
- Ribeiro et al. (2022) L.A. Ribeiro, M.L. Pereira, W.F. Giozza, R.M. Tromer, and Douglas S. Galvão, “Thermal stability and fracture patterns of a recently synthesized monolayer fullerene network: A reactive molecular dynamics study,” Chemical Physics Letters 807, 140075 (2022).
- Zhao et al. (2024) Xiao-Kun Zhao, Yang-Yang Zhang, Jing Zhao, Han-Shi Hu, and Jun Li, “Understanding the electronic structure and chemical bonding in the 2d fullerene monolayer,” Inorg. Chem. 63, 11572–11582 (2024).
- Nagel et al. (1999) P. Nagel, V. Pasler, S. Lebedkin, A. Soldatov, C. Meingast, B. Sundqvist, P.-A. Persson, T. Tanaka, K. Komatsu, S. Buga, and A. Inaba, “C60 one- and two-dimensional polymers, dimers, and hard fullerite: Thermal expansion, anharmonicity, and kinetics of depolymerization,” Phys. Rev. B 60, 16920–16927 (1999).
- Arvanitidis et al. (2003) J. Arvanitidis, Konstantinos Papagelis, Serena Margadonna, Kosmas Prassides, and Andrew N. Fitch, “Temperature-induced valence transition and associated lattice collapse in samarium fulleride,” Nature 425, 599–602 (2003).
- Kwon et al. (2004) Young-Kyun Kwon, Savas Berber, and David Tománek, “Thermal contraction of carbon fullerenes and nanotubes,” Phys. Rev. Lett. 92, 015901 (2004).
- Brown et al. (2006) S. Brown, J. Cao, J. L. Musfeldt, N. Dragoe, F. Cimpoesu, S. Ito, H. Takagi, and R. J. Cross, “Search for microscopic evidence for molecular level negative thermal expansion in fullerenes,” Phys. Rev. B 73, 125446 (2006).
- Huang et al. (2016b) Liang-Feng Huang, Xue-Zeng Lu, Emrys Tennessen, and James M.Rondinelli, “An efficient ab-initio quasiharmonic approach for the thermodynamics of solids,” Computational Materials Science 120, 84–93 (2016b).
- Pallikara and Skelton (2021) Ioanna Pallikara and Jonathan M. Skelton, “Phase stability of the tin monochalcogenides SnS and SnSe: a quasi-harmonic lattice-dynamics study,” Phys. Chem. Chem. Phys. 23, 19219–19236 (2021).
- Gonze (1995) Xavier Gonze, “Adiabatic density-functional perturbation theory,” Phys. Rev. A 52, 1096–1114 (1995).
- Baroni et al. (2001) Stefano Baroni, Stefano de Gironcoli, Andrea Dal Corso, and Paolo Giannozzi, “Phonons and related crystal properties from density-functional perturbation theory,” Rev. Mod. Phys. 73, 515–562 (2001).
- Kresse and Furthmüller (1996a) G. Kresse and J. Furthmüller, “Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set,” Phys. Rev. B 54, 11169–11186 (1996a).
- Kresse and Furthmüller (1996b) G. Kresse and J. Furthmüller, “Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set,” Computational Materials Science 6, 15 – 50 (1996b).
- Dove (1993) Martin T. Dove, Introduction to Lattice Dynamics (Cambridge University Press, 1993).
- Togo et al. (2008) Atsushi Togo, Fumiyasu Oba, and Isao Tanaka, “First-principles calculations of the ferroelastic transition between rutile-type and CaCl2-type SiO2 at high pressures,” Phys. Rev. B 78, 134106 (2008).
- Togo and Tanaka (2015) Atsushi Togo and Isao Tanaka, “First principles phonon calculations in materials science,” Scripta Materialia 108, 1–5 (2015).
- Peng et al. (2019) Bo Peng, Ivona Bravić, Judith L. MacManus-Driscoll, and Bartomeu Monserrat, “Topological semimetallic phase in promoted by temperature,” Phys. Rev. B 100, 161101 (2019).
- Mortazavi (2023) Bohayra Mortazavi, “Structural, electronic, thermal and mechanical properties of c60-based fullerene two-dimensional networks explored by first-principles and machine learning,” Carbon 213, 118293 (2023).
- Kah et al. (2015) C B Kah, M Yu, P Tandy, C S Jayanthi, and S Y Wu, “Low-dimensional boron structures based on icosahedron b12,” Nanotechnology 26, 405701 (2015).
- Sun et al. (2017) Xu Sun, Xiaofei Liu, Jun Yin, Jin Yu, Yao Li, Yang Hang, Xiaocheng Zhou, Maolin Yu, Jidong Li, Guoan Tai, and Wanlin Guo, “Two-dimensional boron crystals: Structural stability, tunable properties, fabrications and applications,” Adv. Funct. Mater. 27, 1603300 (2017).
- Zhang et al. (2017) Zhuhua Zhang, Evgeni S. Penev, and Boris I. Yakobson, “Two-dimensional boron: structures, properties and applications,” Chem. Soc. Rev. 46, 6746–6763 (2017).
- Kroto (1987) H. W. Kroto, “The stability of the fullerenes Cn, with = 24, 28, 32, 36, 50, 60 and 70,” Nature 329, 529–531 (1987).
- Hallett et al. (1995) R. P. Hallett, K. G. McKay, S. P. Balm, A. W. Allaf, H. W. Kroto, and A. J. Stace, “Reaction studies of carbon clusters,” Zeitschrift für Physik D Atoms, Molecules and Clusters 34, 65–70 (1995).
- Xu et al. (2023) Yu-He Xu, Wen-Juan Tian, Alvaro Muñoz-Castro, Gernot Frenking, and Zhong-Ming Sun, “An all-metal fullerene: [k@au12sb20]5-,” Science 382, 840–843 (2023).
- Wu and Peng (2025) Jiaqi Wu and Bo Peng, “Smallest [5,6]fullerene as building blocks for 2d networks with superior stability and enhanced photocatalytic performance,” J. Am. Chem. Soc. 147, 1749–1757 (2025).
- Zhang et al. (2025) Yuxuan Zhang, Yifan Xie, Hao Mei, Hui Yu, Minjuan Li, Zexiang He, Wentao Fan, Panpan Zhang, Antonio Gaetano Ricciardulli, Paolo Samorì, Mengmeng Li, and Sheng Yang, “Electrochemical synthesis of 2d polymeric fullerene for broadband photodetection,” Adv. Mater. 37, 2416741 (2025).
- Wang et al. (2023) Taotao Wang, Li Zhang, Jinbao Wu, Muqing Chen, Shangfeng Yang, Yalin Lu, and Pingwu Du, “Few-layer fullerene network for photocatalytic pure water splitting into h2 and h2o2,” Angew. Chem. Int. Ed. 62, e202311352 (2023).
- Wang et al. (2024) Qingjie Wang, Yongqiang Qiao, Kaiyue Zhao, Peixian Zhang, Huan Zhao, Juan Guo, Xinwei Shi, Erjun Liang, and Qilong Gao, “Zero thermal expansion in kxmnxin2-x(moo4)3 based materials,” Acta Materialia 281, 120358 (2024).
- Yao et al. (2019) Zi-Shuo Yao, Hanxi Guan, Yoshihito Shiota, Chun-Ting He, Xiao-Lei Wang, Shu-Qi Wu, Xiaoyan Zheng, Sheng-Qun Su, Kazunari Yoshizawa, Xueqian Kong, Osamu Sato, and Jun Tao, “Giant anisotropic thermal expansion actuated by thermodynamically assisted reorientation of imidazoliums in a single crystal,” Nature Communications 10, 4805 (2019).
- Bond (2021) Andrew D. Bond, “A survey of thermal expansion coefficients for organic molecular crystals in the Cambridge Structural Database,” Acta Crystallographica Section B 77, 357–364 (2021).
- van der Lee and Dumitrescu (2021) Arie van der Lee and Dan G. Dumitrescu, “Thermal expansion properties of organic crystals: a csd study,” Chem. Sci. 12, 8537–8547 (2021).
- Saha and Veluthaparambath (2024) Binoy K. Saha and Ragima V. P. Veluthaparambath, “Roles of molecular volume, surface area, heteroatoms, and hydrogen bonds on the thermal expansion of organic crystals,” Crystal Growth & Design 24, 3467–3472 (2024).
- Blank et al. (1998) V.D. Blank, S.G. Buga, G.A. Dubitsky, N. R. Serebryanaya, M.Yu. Popov, and B. Sundqvist, “High-pressure polymerized phases of c60,” Carbon 36, 319–343 (1998).
- Masago et al. (2006) Akira Masago, Koun Shirai, and Hiroshi Katayama-Yoshida, “Crystal stability of - and -boron,” Phys. Rev. B 73, 104102 (2006).
Appendix A Thermal expansion of 2D icosahedral B12 networks
To demonstrate the wide applicability of our design principle beyond carbon-based materials, we provide another example based on the icosahedral B12 units in Fig. 5(a), as these clusters form the most stable allotropes of boron including the and phases Masago et al. (2006). We explicitly compute thermal expansion of monolayer qHP B12 networks and find the same trend in Fig. 5(b).